# DPMambaIR: All-in-One Image Restoration via Degradation-Aware Prompt State Space Model

### v1

@@ Abstract.
All-in-One image restoration aims to address multiple image degradation problems using a single model, significantly reducing training costs and deployment complexity compared to traditional methods that design dedicated models for each degradation type. Existing approaches typically rely on Degradation-specific models or coarse-grained degradation prompts to guide image restoration. However, they lack fine-grained modeling of degradation information and face limitations in balancing multi-task conflicts. To overcome these limitations, we propose DPMambaIR, a novel All-in-One image restoration framework. By integrating a Degradation-Aware Prompt State Space Model (DP-SSM) and a High-Frequency Enhancement Block (HEB), DPMambaIR enables fine-grained modeling of complex degradation information and efficient global integration, while mitigating the loss of high-frequency details caused by task competition. Specifically, the DP-SSM utilizes a pre-trained degradation extractor to capture fine-grained degradation features and dynamically incorporates them into the state space modeling process, enhancing the model’s adaptability to diverse degradation types. Concurrently, the HEB supplements high-frequency information, effectively addressing the loss of critical details, such as edges and textures, in multi-task image restoration scenarios. Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, with 27.69dB and 0.893 in PSNR and SSIM, respectively. These results highlight the potential and superiority of DPMambaIR as a unified solution for All-in-One image restoration.
All-in-One Image Restoration, State Space Model, Prompt Learning
1. Introduction
@@
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Recent advances in deep learning have significantly improved performance in Degradation-specific restoration problems, such as image denoising (Kim et al., 2024; Zeng et al., 2024; Ryou et al., 2024), deblurring (Chen et al., 2024b; Wu et al., 2024a; Mao et al., 2024), deraining(Chen et al., 2024a, 2023; Ba et al., 2022; Peng et al., 2020, 2021; He et al., 2024b), dehazing(Zhang et al., 2024e; Wang et al., 2024c; Zheng et al., 2023; Wu et al., 2023) and low-light enhancement(Ma et al., 2022; Cai et al., 2023; Hai et al., 2023; Wang et al., 2022, 2023b). However, traditional methods typically rely on Degradation-specific models explicitly tailored to individual degradation types. These approaches often require prior knowledge of the degradation type, which limits their practicality in real-world scenarios, such as autonomous driving and nighttime surveillance, where diverse and unknown degradations frequently coexist.
@@
Figure 1. Comparison with SOTA All-in-One Image Restoration Methods on mixed Dataset. Our DPMambaIR achieves consistently better performance.
@@
To address this limitation, recent studies explore All-in-One image restoration frameworks that unify multiple restoration tasks into a single model. Existing methods can be broadly categorized into Mixture-of-Experts (MoE)-based and Prompt-based approaches. MoE-based methods, such as MoFME (Zhang et al., 2024b), dynamically route tasks to specialized experts using an uncertainty-aware router, enabling effective learning of diverse expert functions. However, these methods suffer from high computational costs, increased inference complexity as the number of experts grows, and lack explicit degradation modeling, limiting adaptability to unseen or complex degradations. Prompt-based methods, like OneRestore (Guo et al., 2024a), guide restoration using degradation prompts, integrating scene descriptors via cross-attention mechanisms for flexible and adaptive strategies. While effective, these methods rely on coarse prompts that fail to capture nuanced degradation characteristics and treat prompts as static inputs, limiting their ability to adapt to varying degradation intensities or specific requirements, leading to suboptimal results.
@@
While these approaches reduce computational cost and deployment complexity by leveraging feature sharing and multi-task learning, two key challenges remain unresolved. First, existing models lack fine-grained degradation awareness and often rely on coarse degradation labels to guide the restoration process. For example, images with varying degrees of haze are typically assigned a single haze label, failing to capture the nuanced variations in degradation levels. This simplification leads to suboptimal restoration performance, as the model cannot adapt to the diverse characteristics of different degradations. Second, current methods struggle to balance the competing requirements of varying restoration tasks. For instance, training low-light enhancement, which focuses on restoring low-frequency contrast information, alongside deblurring, which prioritizes the recovery of fine local details, often results in task competition. This inherent task competition makes it challenging to optimize a single network for all restoration tasks simultaneously and remains a fundamental challenge for All-in-One image restoration frameworks.
@@
To address these challenges, we propose a novel All-in-One image restoration framework, DPMambaIR. The framework is built upon a Degradation-aware Prompt State Space Model (DP-SSM) and a lightweight High-frequency Enhancement Block (HEB). Specifically, DP-SSM employs a pre-trained degradation extractor to extract degradation embeddings, which are used to modulate the state transition, input, and output matrices of the state space equations. This enables fine-grained degradation modeling and efficient global degradation integration. Meanwhile, HEB mitigates the loss of high-frequency details caused by task competition among different restoration tasks. Extensive experiments on a mixed dataset containing seven degradation types demonstrate the superiority of DPMambaIR. As shown in Figure 1, DPMambaIR achieves state-of-the-art performance in terms of PSNR and SSIM, outperforming existing All-in-One methods such as AdaIR and OneRestore. Moreover, DPMambaIR achieves competitive results across individual tasks, including deraining, low-light enhancement, deblurring, and dehazing, showcasing its potential as an effective and robust All-in-One image restoration solution.
The contributions of this paper can be summarized:
@@
    We propose a Degradation-aware Prompt State Space Model (DP-SSM) that incorporates fine-grained degradation features into the state space modeling process, enabling both global degradation modeling and fine-grained degradation awareness to effectively handle diverse degradation types within a unified framework.
@@
Figure 2. The overall framework of DPMambaIR consists of a shallow feature extractor and multiple DPSS blocks, with an upsampler implemented by pixel shuffle and a downsampler by pixel unshuffle. Additionally, a degradation extractor is included to extract degradation embeddings, which are then inserted into the DPSS blocks.
2. Related Work
2.1. All-in-One Image Restoration
@@
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Traditional methods, such as Dark Channel Prior (He et al., 2010) and Color Line Prior (Fattal, 2014), rely on handcrafted priors to model degradation but struggle with complex scenarios due to predefined assumptions. The rise of deep learning (LeCun et al., 2015) has significantly advanced tasks like such as image super-resolution(Peng et al., 2024c, d, 2025c; Wang et al., 2024b; Saharia et al., 2022; Gao et al., 2023), image denoising (Kim et al., 2024; Zeng et al., 2024; Ryou et al., 2024), deblurring (Chen et al., 2024b; Wu et al., 2024a; Mao et al., 2024), deraining(Chen et al., 2024a; Pan et al., 2023; Chen et al., 2023; Ba et al., 2022; Peng et al., 2020, 2021; He et al., 2024b), dehazing(Zhang et al., 2024e; Wang et al., 2024c; Zheng et al., 2023; Wu et al., 2023), desnowing(Guo et al., 2025; Chen et al., 2021; Zhang et al., 2021) and low-light enhancement(Ma et al., 2022; Cai et al., 2023; Hai et al., 2023; Wang et al., 2022, 2023b; Ma et al., 2024). These approaches, leveraging convolutional neural networks (CNN)-based (LeCun et al., 1989), transformer-based (Vaswani et al., 2017; Pan et al., 2025a, 2024), or diffusion-based (Ho et al., 2020) architectures, enable powerful feature learning for image restoration. However, most existing works focus on addressing a single type of degradation, which limits their applicability in real-world scenarios where images often suffer from diverse and unknown types of degradation.
@@
Recent advancements in All-in-One image restoration frameworks focus on enhancing flexibility and efficiency by leveraging various multi-task learning techniques such as Mixture-of-Experts (MoE) and text-based or vision-based prompts to handle diverse restoration tasks. For example, AirNet(Li et al., 2022) employs contrastive learning to extract degradation representations for guiding restoration. PromptIR(Potlapalli et al., 2023) integrates end-to-end prompt generation and interaction to adaptively handle different degradation types, while OneRestore(Guo et al., 2024a) aligns text-based degradation categories with visual features as prompts. IDR(Zhang et al., 2023a) uses meta-learning to decompose degradations, enabling generalization across diverse scenarios. MoFME(Zhang et al., 2024b), on the other hand, employs an uncertainty-aware router to dynamically assign Degradation-specific features to expert modules, allowing it to specialize in multiple tasks. Similarly, AdaIR(Cui et al., 2025) achieves adaptive restoration by decoupling degradations in the frequency domain. Additionally, methods like DA-CLIP(Luo et al., 2023a) leverage the pre-trained vision-language model (VLM) for prompt extraction, but the significant computational overhead introduced by VLM limits their practical efficiency. However, these methods either limit the representation of degradation information to coarse-grained degradation categories or lack a clear mechanism for extracting degradation features, which often results in suboptimal performance. In contrast, we propose a novel All-in-One image restoration framework, DPMambaIR, which is capable of extracting fine-grained degradation information and incorporating a state-space modeling process, effectively addressing diverse degradation challenges in image restoration.
2.2. State Space Model
State Space Model (SSM), originally introduced in the 1960s for control systems(Kalman, 1960), has recently gained significant attention in computer vision due to its ability to model long-range dependencies with linear computational complexity. The first effort to incorporate SSM into vision tasks was initiated by Vision Mamba(Zhu et al., 2024), which proposed the Vision State Space Module (VSSM) to extend SSM to visual applications. Vision Mamba demonstrated superior performance compared to Vision Transformers (ViTs)(Dosovitskiy et al., 2020), while maintaining lower model complexity, thereby laying a solid foundation for subsequent advancements in the field of computer vision. The success of VSSM has inspired extensive research into the application of SSM across various vision tasks, such as object detection(Zhang et al., 2024a; Yang et al., 2025; Liu et al., 2024; Li et al., 2024c; Zhang et al., 2023b, c), semantic segmentation(Xiao et al., 2024; Huang et al., 2023; Zhang et al., 2024c), and image classification(Zhang et al., 2024d; Li et al., 2024b; Yao et al., 2024; jc5sheng2024dualmamba). Among these, MambaIR(Guo et al., 2024c) is the first approach to apply SSM to image restoration. By leveraging SSM, MambaIR achieves enhanced computational efficiency and improved global feature modeling capabilities, resulting in significant improvements in perceptual quality for restoration tasks. Furthermore, other SSM-based approaches have also emerged for various image restoration tasks(Hu and Yin, 2024; Li et al., 2024d; Guo et al., 2024b; He et al., 2024c; Sun et al., 2024b; Li et al., 2024a; Wu et al., 2024c; Xia et al., 2024; Peng et al., 2025a; Di et al., 2024; Peng et al., 2025b; Pan et al., 2025b; Qi et al., 2025; Sun et al., 2024a; Wu et al., 2024b; He et al., 2024a; Du et al., 2024; He et al., 2024d; Jiang et al., 2024; Wang et al., 2024a; Li et al., 2023; Conde et al., 2024; Ren et al., 2024; Yan et al., 2023), showcasing the strong potential and promising opportunities of SSM in this domain. However, while SSM excels at modeling long-range interactions, they inherently lack a degradation-aware mechanism, which is crucial for tasks that involve handling diverse degradations. This limitation makes it difficult to balance differences between multiple tasks during modeling. This limitation hinders the performance of the model in All-in-One image restoration tasks.
3. Method
3.1. Problem Definition
All-in-One Image Restoration aims to address a wide range of image degradation types within a unified framework. Specifically, given a degraded image $\boldsymbol{I}_{D}\in\mathbb{R}^{3\times H\times W}$, where $D$ denotes the unknown degradation process, our goal is to restore the clean image $\boldsymbol{\hat{O}}\in\mathbb{R}^{3\times H\times W}$ using a unified model $\mathcal{F}$. This can be formulated as:
(1) $\boldsymbol{\hat{O}}=\mathcal{F}(\boldsymbol{I}_{D}).$
where $f$ is designed to generalize across various degradation types, such as noise, blur, compression artifacts, and weather effects, without explicit knowledge of $D$. The degradation process $D$ can be mathematically represented as:
(2) $\boldsymbol{I}_{D}=D(\boldsymbol{I}_{C})+\boldsymbol{\eta}.$
Here, $\boldsymbol{I}_{C}\in\mathbb{R}^{3\times H\times W}$ denotes the original clean image, $D$ represents the degradation operator (e.g., blur kernel, down-sampling, or quality compression), and $\boldsymbol{\eta}$ signifies additive noise. The main challenge arises from the fact that $D$ is usually unknown and varies across tasks, making it difficult to design a universal model capable of effectively handling all possible degradations.
3.2. Traditional State Space Model
@@
In Image Restoration, Convolutional Neural Networks (CNN)(LeCun et al., 2015) and Transformer(Vaswani et al., 2017) have been widely adopted due to their strong feature extraction and representation learning capabilities. CNN excels at capturing local patterns, making them suitable for tasks requiring high spatial resolution and fine details. However, their limited receptive field often hinders their ability to model long-range dependencies, which are critical for handling global degradations. On the other hand, Transformer leverages self-attention mechanisms to capture global context effectively. Despite their success, Transformer suffers from quadratic complexity with respect to the input size, leading to high computational costs and memory consumption, especially for high-resolution images. To address these limitations, researchers have turned to Mamba, which leverages State Space Model (SSM) to achieve linear complexity and long-range dependency modeling, making it well-suited for Image Restoration tasks.
@@
Let’s briefly review the traditional State Space Model (SSM), which maps sequence input $x(t)$ to output $y(t)$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$ and can be represented as a linear ordinary differential equation:
(3) $\dot{h}(t)$ $=Ah(t)+Bx(t),$
@@
where $N$ is the state size, $A\in\mathbb{R}^{N\times N}$, $B\in\mathbb{R}^{N\times 1}$, $C\in\mathbb{R}^{1\times N}$, and $D\in\mathbb{R}$. Discretized using a zero-order hold as:
(4) $\overline{A}$ $=\exp(\Delta A),$
$\overline{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B.$
After the discretization, the discretized version of Eq. (3) with step size $\Delta$ can be rewritten as:
(5) $h_{i}$ $=\overline{A}h_{i-1}+\overline{B}x_{i},$
@@
The advent of State Space Model has garnered significant attention for its ability to perform global modeling with linear complexity in image restoration. Existing Mamba-based methods(Zhu et al., 2024; Guo et al., 2024c, b) typically adopt specific scanning strategies to unfold 2D images into 1D sequences, enabling sequential causal modeling using Mamba. However, in All-in-One image restoration tasks, traditional SSM still faces two key limitations: (a) Limited Global Context Utilization: During the sequential causal modeling process, the $i$-th pixel can only access the previous $i-1$ pixels in the sequence, preventing effective utilization of global contextual information. To address this limitation, existing methods often employ parallel multi-directional scanning to expand the receptive field. However, this inevitably increases computational complexity. Furthermore, experiments in existing work(Guo et al., 2024c) reveal that parallel multi-directional scanning contributes minimally to performance improvements in low-level vision tasks. (b) Lack of Effective Degradation-Aware Mechanisms: Traditional SSM lack mechanisms to effectively perceive degradation information, making it difficult to balance differences between multiple tasks during modeling. This limitation hinders the performance of the model in All-in-One image restoration tasks.
3.3. DPMambaIR

### v2

@@ Abstract
All-in-One image restoration aims to address multiple image degradation problems using a single model, offering a more practical and versatile solution compared to designing dedicated models for each degradation type. Existing approaches typically rely on Degradation-specific models or coarse-grained degradation prompts to guide image restoration. However, they lack fine-grained modeling of degradation information and face limitations in balancing multi-task conflicts. To overcome these limitations, we propose DPMambaIR, a novel All-in-One image restoration framework that introduces a fine-grained degradation extractor and a Degradation-Aware Prompt State Space Model (DP-SSM). The DP-SSM leverages the fine-grained degradation features captured by the extractor as dynamic prompts, which are then incorporated into the state space modeling process. This enhances the model’s adaptability to diverse degradation types, while a complementary High-Frequency Enhancement Block (HEB) recovers local high-frequency details. Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, with 27.69dB and 0.893 in PSNR and SSIM, respectively. These results highlight the potential and superiority of DPMambaIR as a unified solution for All-in-One image restoration.
I Introduction
@@
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Recent advances in deep learning have significantly improved performance in Degradation-specific restoration tasks, such as image denoising [1, 2], deblurring [3], deraining[4], dehazing[5, 6] and low-light enhancement[7]. However, traditional methods typically rely on Degradation-specific models explicitly tailored to individual degradation types. These approaches often require prior knowledge of the degradation type, which limits their practicality in real-world scenarios, such as autonomous driving and nighttime surveillance, where diverse and unknown degradations frequently coexist.
@@
Figure 1: Comparison with SOTA All-in-One Image Restoration Methods on mixed Dataset. Our DPMambaIR achieves consistently better performance.
@@
To address this limitation, recent studies explore All-in-One image restoration frameworks that unify multiple restoration tasks into a single model. Existing methods can be broadly categorized into Mixture-of-Experts (MoE)-based and Prompt-based approaches. MoE-based methods, such as MoFME [8], dynamically route tasks to specialized experts using an uncertainty-aware router, enabling effective learning of diverse expert functions. However, these methods suffer from high computational costs, increased inference complexity as the number of experts grows, and lack explicit degradation modeling, limiting adaptability to unseen or complex degradations. Prompt-based methods, like OneRestore [9], guide restoration using degradation prompts, integrating scene descriptors via cross-attention mechanisms for flexible and adaptive strategies. While effective, these methods rely on coarse prompts that fail to capture nuanced degradation characteristics and treat prompts as static inputs, limiting their ability to adapt to varying degradation intensities or specific requirements, leading to suboptimal results.
@@
While these methods have demonstrated success in learning multiple tasks for All-in-One image restoration, two key challenges remain unresolved. First, existing models lack a degradation extractor capable of capturing fine-grained degradation features and often rely on coarse degradation labels to guide the restoration process. For example, images with varying degrees of haze are typically assigned a single haze label, failing to capture the nuanced variations in degradation levels. This simplification leads to suboptimal restoration performance, as the model cannot adapt to the diverse characteristics of different degradations. Second, current methods struggle to balance the competing requirements of varying restoration tasks. For instance, training low-light enhancement, which focuses on restoring low-frequency contrast information, alongside deblurring, which prioritizes the recovery of fine local details. This inherent task competition makes it challenging to optimize a single network for all restoration tasks simultaneously and remains a fundamental challenge for All-in-One image restoration frameworks.
To address these challenges, we propose DPMambaIR, a novel All-in-One image restoration framework built upon a regression-based degradation extractor and a Degradation-Aware Prompt State Space Model (DP-SSM). Unlike prior approaches that use static or class-based prompts, our method employs a degradation extractor to generate prompt vectors that encode both semantic categories and degradation distribution characteristics. These prompts are then used to dynamically modulate the state transition, input, and output matrices of a state space model, enabling fine-grained degradation modeling and adaptive coordination across frequency bands. This mechanism allows our model to better manage task competition and enhance restoration quality across diverse degradations. Moreover, while State Space Model such as Mamba exhibit strong global modeling capabilities, they are inherently biased toward low-frequency restoration [10]. To better recover local high-frequency details, we introduce a lightweight High-Frequency Enhancement Block (HEB) as a supplementary component, which delivers performance gains while incurring only negligible computational overhead.
@@
Extensive experiments on a mixed dataset containing seven degradation types demonstrate the superiority of DPMambaIR. As shown in Fig. 1, DPMambaIR achieves state-of-the-art performance in terms of PSNR and SSIM, outperforming existing All-in-One methods such as AdaIR and OneRestore. Moreover, DPMambaIR achieves competitive results across individual tasks, including deraining, low-light enhancement, deblurring, and dehazing, showcasing its potential as an effective and robust All-in-One image restoration solution.
The contributions of this paper are summarized as follows:
@@
    We design a Degradation-aware Prompt State Space Model (DP-SSM), which enhances the long-range modeling capability of Mamba by integrating degradation priors at multiple structural levels.
@@
II Related Work
II-A All-in-One Image Restoration
@@
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Traditional methods, such as Dark Channel Prior [11] and Color Line Prior [12], rely on handcrafted priors to model degradation but struggle with complex scenarios due to predefined assumptions. The rise of deep learning [13] has significantly advanced tasks like such as image super-resolution[14, 15, 16], image denoising [1, 2], deblurring [3], deraining[4], dehazing[5], desnowing[17] and low-light enhancement[7]. These approaches, leveraging convolutional neural networks (CNN)-based [18], transformer-based [19], or diffusion-based [20] architectures, enable powerful feature learning for image restoration. However, most existing works focus on addressing a single type of degradation, which limits their applicability in real-world scenarios where images often suffer from diverse and unknown types of degradation.
@@
Recent advancements in All-in-One image restoration frameworks focus on enhancing flexibility and efficiency by leveraging various multi-task learning techniques such as Mixture-of-Experts (MoE) and text-based or vision-based prompts to handle diverse restoration tasks. For example, AirNet[21] employs contrastive learning to extract degradation representations for guiding restoration. PromptIR[22] integrates end-to-end prompt generation and interaction to adaptively handle different degradation types, while OneRestore[9] aligns text-based degradation categories with visual features as prompts. IDR[23] uses meta-learning to decompose degradations, enabling generalization across diverse scenarios. MoFME[8], on the other hand, employs an uncertainty-aware router to dynamically assign Degradation-specific features to expert modules, allowing it to specialize in multiple tasks. Similarly, AdaIR[24] achieves adaptive restoration by decoupling degradations in the frequency domain. Additionally, methods like DA-CLIP[25] leverage the pre-trained vision-language model (VLM) for prompt extraction, but the significant computational overhead introduced by VLM limits their practical efficiency.
Despite their effectiveness, most prompt-based methods rely on coarse-grained class labels or static degradation tokens, which limits their ability to model continuous degradation distributions or adapt to image-specific degradation characteristics. In contrast, our method extracts fine-grained, regression-level degradation embeddings that encode both semantic and statistical degradation features. These embeddings are used to dynamically modulate the state-space parameters of the restoration backbone, enabling degradation-aware processing throughout the network. This facilitates better coordination across restoration tasks and reduces competition between frequency components.
Figure 2: The overall framework of DPMambaIR consists of a shallow feature extractor and multiple DPSS blocks, with an upsampler implemented by pixel shuffle and a downsampler by pixel unshuffle. Additionally, a degradation extractor is included to extract degradation embeddings, which are then inserted into the DPSS blocks.
II-B State Space Model based Image Restoration
@@
In Image Restoration, Convolutional Neural Networks (CNN) [13] and Transformer [19] have been widely adopted due to their strong feature extraction and representation learning capabilities. CNN excels at capturing local patterns, making them suitable for tasks requiring high spatial resolution and fine details. However, their limited receptive field often hinders their ability to model long-range dependencies, which are critical for handling global degradations. On the other hand, Transformer leverages self-attention mechanisms to capture global context effectively. Despite their success, Transformers suffer from quadratic complexity relative to the input size, making them computationally and memory-intensive.
@@
To address these limitations, researchers have turned to the State Space Model (SSM)[26], which provides an efficient framework for modeling long-range dependencies with linear computational complexity. Formally, SSM maps an input sequence $x(t)$ to an output sequence $y(t)$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$, represented by a linear ordinary differential equation:
$h^{\prime}(t)$ $=Ah(t)+Bx(t),$ (1)
@@
where $A\in\mathbb{R}^{N\times N}$, $B\in\mathbb{R}^{N\times 1}$, $C\in\mathbb{R}^{1\times N}$, and $D\in\mathbb{R}$. After discretization with step size $\Delta$, the dynamic equation can be expressed as:
$h_{i}$ $=\overline{A}h_{i-1}+\overline{B}x_{i},$ (2)
@@
where $\overline{A}=\exp(\Delta A)$ and $\overline{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B$.
The introduction of Vision Mamba[27] marked the first effort to incorporate SSM into vision tasks. By proposing the Vision State Space Module (VSSM), it extended SSM to visual applications and demonstrated superior performance over Vision Transformer (ViT)[28], while maintaining lower complexity. This success inspired extensive research into SSM across diverse tasks such as object detection[29], semantic segmentation[30], and image classification[31].
In the context of Image Restoration, MambaIR[32] pioneered the integration of SSM, leveraging its linear complexity and long-range modeling capability to achieve significant improvements in both computational efficiency and perceptual quality. This was followed by several variants targeting specific restoration tasks[33, 34, 35], collectively showcasing the strong potential of SSM in this domain. Existing approaches typically unfold 2D images into 1D sequences and apply sequential causal modeling. However, this design still faces two limitations: (a) Limited Global Context Utilization: The $i$-th pixel can only access the previous i1i-1 pixels in the sequence, restricting the effective use of global context. Parallel multi-directional scanning has been explored to enlarge the receptive field, but it increases computational cost while providing limited gains in low-level vision tasks[32]. (b) Lack of Effective Degradation-Aware Mechanisms: Current SSM-based methods primarily target single-task restoration. Their transition and output mappings are fixed or solely input-driven, without leveraging external degradation priors. As a result, they cannot dynamically adapt to diverse degradation types or levels, leading to suboptimal performance in All-in-One image restoration task.
These limitations motivate the introduction of a degradation-aware modulation mechanism into the SSM pipeline. In this work, we propose a Degradation-aware Prompt State Space Model (DP-SSM), where degradation embeddings dynamically modulate the state-space transition and output matrices. This design enables a fine-grained, task-aware mechanism that balances global modeling with degradation-specific guidance, enhancing the flexibility and effectiveness of SSM for complex restoration scenarios.
III Method
III-A Problem Definition
All-in-One Image Restoration aims to address a wide range of image degradation types within a unified framework. Specifically, given a degraded image $\boldsymbol{I}_{D}\in\mathbb{R}^{3\times H\times W}$, where $D$ denotes the degradation process, our goal is to restore the clean image $\boldsymbol{\hat{O}}\in\mathbb{R}^{3\times H\times W}$ using a unified model $\mathcal{F}$. This can be formulated as:
$\boldsymbol{\hat{O}}=\mathcal{F}(\boldsymbol{I}_{D}).$ (3)
where $\mathcal{F}$ is designed to generalize across various degradation types, such as noise, blur, compression artifacts, and weather effects, without explicit knowledge of $Deg$. The degradation process $Deg$ can be mathematically represented as:
$\boldsymbol{I}_{D}=Deg(\boldsymbol{I}_{C})+\boldsymbol{\eta}.$ (4)
Here, $\boldsymbol{I}_{C}\in\mathbb{R}^{3\times H\times W}$ denotes the original clean image, $Deg$ represents the degradation operator (e.g., blur kernel, brightness, or quality compression), and $\boldsymbol{\eta}$ signifies additive noise. The main challenge arises from the fact that $Deg$ is usually unknown and varies across tasks, making it difficult to design a universal model capable of effectively handling all possible degradations.
Unlike degradation-specific restoration, the All-in-One setting requires the model to generalize across a broad spectrum of degradation types without prior knowledge of degradation parameters. This introduces two main challenges: (a) how to represent fine-grained degradation information in a unified way; and (b) how to dynamically adapt the model behavior based on such information during inference.
Figure 3: (a) Real-World Image Degradation Modeling Process. (b) Architecture details of the Large version degradation extractor and reconstructor. (c) Architecture details of the normal version degradation extractor and reconstructor.
III-B DPMambaIR